OpenStudy (kitkat16):

Please help!!! What is the sum of the first eight terms of a geometric series whose first term is 3 and whose common ratio is 1/2?

OpenStudy (kitkat16):

error... the common ratio is 1/2

OpenStudy (kitkat16):

Is it 5.9765625=765/128 ???

OpenStudy (kitkat16):

@ganeshie8 can you check please

OpenStudy (kitkat16):

@hartnn can you help me please!!! is this correct 5.9765625=765/128

OpenStudy (kitkat16):

so the common ratio is 1/2 yes?

OpenStudy (kitkat16):

the term is 3

hartnn (hartnn):

yes, r =1/2 same formula \(S_n = a_1\dfrac{r^n-1}{r-1}\) a1 = 1st term = 3 n= 8

OpenStudy (kitkat16):

0.5+0.25+0.125

hartnn (hartnn):

765/128 is correct:)

OpenStudy (kitkat16):

thank you